OpenStudy (anonymous):

Simplify 3 sqrt 7 divided by 5 sqrt 7. Hoping I set this up right

Nnesha (nnesha):

\[\huge\rm \frac{ 3\sqrt{7} }{ 5\sqrt{7} }\] like this ?

OpenStudy (anonymous):

Yes I wasn't sure I was writing it correctly

Nnesha (nnesha):

can you cancel out anything ?

OpenStudy (anonymous):

I would say the 7's

Nnesha (nnesha):

yes right sqrt{7 } that's it

OpenStudy (anonymous):

Ok so would that be 7 1/15?

Nnesha (nnesha):

hmm how ? how did you get 15

OpenStudy (anonymous):

No I mean 7 1/5

OpenStudy (anonymous):

Hi Nnesha was my answer correct?

jimthompson5910 (jim_thompson5910):

The square root of 7 terms cancel \[\Large \frac{ 3\sqrt{7} }{ 5\sqrt{7} }\] \[\Large \frac{ 3\color{red}{\sqrt{7}} }{ 5\color{red}{\sqrt{7}} }\] \[\Large \frac{ 3\cancel{\color{red}{\sqrt{7}}} }{ 5\cancel{\color{red}{\sqrt{7}}} }\] \[\Large \frac{3}{5}\] Just leaving the fraction 3/5. I'm not sure how you got 7&1/5

OpenStudy (anonymous):

Ok my choices for answers were 7 1/5 7 5/3 7 2/15 That's what threw me off. I had the canceling done correctly but wasn't sure how these were the choices. But thank you for assisting me I appreciate it😀

OpenStudy (triciaal):